Project

General

Profile

Ops/Comm Scheduling meeting : 18/7/2017

Location : CR21 at 14:00 pm

Phone access : New phonebridge number
UK: 08082380274
US: 631.621.5253
Access code: 1779339517

  • Actions
    • Hall Manager - water system change-over procedure (stands) - have prescription and draft document (Colin/Josef), stands
    • Check paper logbook for issues addressed there - stands
  • DC update
  • State of air conditioning
  • Hydrogen system update and plans
  • Cycle 2017 Issues
  • Update on Financial Restrictions in Technology and Engineering Subjects (CWhyte)
  • AOB
    • DS controls (LHe level fluctuations) and State Machine control

Minutes

  • Actions
    • Hall Manager - water system change-over procedure (stands) - have laminated document ready to post (Colin/Josef), need more visible labelling of valves; stands
    • Check paper logbook for issues addressed there (SBoyd) - stands
  • DC update
  • State of air conditioning
    The aircon company have been in. Unit 2 (defrosted) has been checked over and shows no errors; we can swap back to using it after we know what's been checked.
    The FC showed a pressure rise after unit 5 was turned off; or is this a diurnal variation?
    The fix for the units 2 and 3 swap is in the code but not active; needs the IOC restarting. AKurup will do this afternoon.
  • Hydrogen system update and plans (JBoehm)
    The FC is back in and the large leak has gone away (swarf across end-flange O-ring); final leak checking in progress. All the temperature readings have come back OK. If the vacuum is good then can start Neon test today/tomorrow.
    Eamonn is making pressure pads to stop the Fischer connectors rotating upon disconnection/reconnection.
    LH2 safety tour at 9 a.m. Thursday.
    JB is working on paperwork. Has made a short version of the instructions - test with Craig.
    The GLIMOS needs to know where the paperwork is. It should be on the wiki - but where on the wiki? On the LH2 wiki but linked to from the Ops Subsystem documentation section.
    The lift-off plates are in R9; preparing for test today or tomorrow.
    Steve Harrison analysis suggests a single 100 mm duct is sufficient.
    Need to think about the integration between LH2 and the rest of MICE. (e.g.) Is the requirement for locking Hall and Controlled Entries necessary? JBoehm/CMacwaters/HNebrensky to discuss by Thursday
  • Cycle 2017 Issues
    The list has been reviewed by SBoyd and CWhyte. The meeting went through the C&M and Beamline sections of the list, skipping over "Other":
    • C&M: Long-term alarms should be individually Disabled, not just silenced. We will need to restart the Alarm Handler regularly to get rid of anything inappropriately disabled, but will then need to work through the resulting alarms to identify whether they are known issues or not. This for the MOM to oversee. We should do at 9am each Wednesday; add to the Run Meeting agenda as reminder. Need to keep the list of known alarms somehow.
      The Tracker vacuum gauges should now be reading out.
      Cryo-compressor communications: the cable is OK, the comms has been reset, believed fixed.
      SecurityProbe PVs going stale: stands. Can we check the age of the PV in Nagios? (Paolo)
      VME PCs: Remove the CPU temperatures as the reading is bogus. AKurup has checked with Pete Owens: the system temperature limits seem reasonable and thought that clogged fans might be the problem. But all are supposed to have been cleaned out - action on the DL electrical list.
    • Beamline: HNebrensky to show CWhyte the Beamstop issues as no John. Circuit 1 not urgent; restraint frame is.
      The new target PC is now running - at last - but the R78 test rig has stopped with a BPS error; HNebrensky will go after meeting and tweak some parameters at Sheffield's request.
  • Update on Financial Restrictions in Technology and Engineering Subjects (CWhyte)
    Yes!
  • AOB
    • DS controls (LHe level fluctuations) and State Machine control
      The State Machine being in an inappropriate state caused the scan rate of a particular PV to be reduced, such that the filtering applied to said PV smoothed out real, rapid changes in LHe level, effectively blinding the control loop to the variation hence it didn't respond.
      DS has been left in manual control. Mark did try changing the DS' internal state in the hope it would kick the SM into the correct state, but instead tripped off the fridge and we lost gas. Neither Mark, Henry or Craig (DC) has access to the MOM State Machine control panel.
      Part of the PV issue is covered by an existing bug fix that wasn't however active on the machine in use - needs a restart of the IOC. AKurup will do this afternoon. Also need to update and reload spreadsheet.
      Outside of running, access to the State Machine control panel is through the DC. Is the password in the document in the MLCR binder?
    • No update regarding John Govans, must plan/proceed on the basis of an indefinite absence
    • FC power supply: Craig talked to Trevor last week, still no progress from Thurlby-Thandar